The position of a particle moving along X axis varies with time as X=10-2t
1
Expert's answer
2021-09-13T12:16:02-0400

Given that ;

"x=10t-2t^2"

So;"V=\frac {dx} {dt} =10-4t"

The particle will come to rest if v = 0

"10-4t=0"

"t=2.5s"